OpenStudy (anonymous):

Hexagon DEFGHI is translated on the coordinate plane below to create hexagon D'E'F'G'H'I': Hexagon DEFGHI and Hexagon D prime E prime F prime G prime H prime I prime on the coordinate plane with ordered pairs at D 2, 5, at E 5, 5, at F 6, 3, at G 5, 1, at H 2, 1, at I 1, 3, at D prime negative 6, negative 2, at E prime negative 3, negative 2, at F prime negative 2, negative 4, at G prime negative 3, negative 6, at H prime negative 6, negative 6, at I prime negative 7, negative 4 Which rule represents the translation of hexagon DEFGHI to hexagon D'E'F'G'H'I'?

OpenStudy (alexandervonhumboldt2):

D(2, 5) => D'(-6, -2) E(5, 5) => E'(-3, -2) F(6, 3) => F'(-2, -4) to find the rule just subtract corrdinate 1 of any pair of points from coordinate 2: -2-6=-8 -4-3=-7 so the rule is (x-8 , y-7)

OpenStudy (anonymous):

Defiantly (x,y)-> (x-8, y-7)